OpenStudy (anonymous):

can anyone deduce sequence definitions from a given sequence (the first six terms)? -1, -3, 1, 17, 51, 109.

OpenStudy (anonymous):

using this might help 􀀀1 = c3 + c2 + c1 + c0 (1) 􀀀3 = 8c3 + 4c2 + 2c1 + c0 (2) 1 = 27c3 + 9c2 + 3c1 + c0 (3) 17 = 64c3 + 16c2 + 4c1 + c0 (4) boxes are - From equation (1) we have c0 = 􀀀1 􀀀 c3 􀀀 c2 􀀀 c1. Substituting this into equations (2), (3) and (4) we obtain

OpenStudy (anonymous):

so i have this solution but i dont really understand it

OpenStudy (kfujioka):

oh yeah, that makes sense.

OpenStudy (anonymous):

it does?

OpenStudy (anonymous):

awwww I was doing it right I just messed it up!

OpenStudy (kfujioka):

You use the finite differences to get a feel for the degree of the polynomial. Since the 3rd finite difference remains constant, that's like saying that the 3rd derivative of the polynomial is constant. Which means a polynomial of at most degree 3. Then you set up a system of equations to solve for the correct coefficients.

OpenStudy (anonymous):

ok the last bit is the troubling stuff (the basics) ie: Then you set up a system of equations to solve for the correct coefficients.

OpenStudy (kfujioka):

Ok, do you get that there are four terms to solve for?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i drew you a picture of what i did... but it looks better on paper.

OpenStudy (kfujioka):

So the general polynomial: \[c_3x ^{3} + c_2x^{2} + c_1x + c_0\] can be equated to the individual terms of your sequence. For the first term plug in x=1, for the second x=2, etc. With 4 unknowns (the c's) you need 4 equations.

OpenStudy (anonymous):

whoops you have to save that, give me a minute to convert it.

OpenStudy (anonymous):

ok i see your still here eliza, so can you explain this to me please (the algebra - seeing as im terribad at it) \[7c _{3} + 3C _{2} + C _{1} = -2\] can you explain how im meant to rearrange for c1?

OpenStudy (anonymous):

\[C _{1}=-7_{c3}-3C _{2}-2\] and i am always here paul o.0

OpenStudy (kfujioka):

subtract \[7c_3 + 3c_2\] from both sides to get \[c_1 = -2 -7c_3-3c_2\]

OpenStudy (anonymous):

oh, its that easy im doing that too much get stuck on the easiest stuff every time...

OpenStudy (kfujioka):

No worries. We all do.

OpenStudy (anonymous):

now i just have to understand the rest of the question...

OpenStudy (anonymous):

with 12c3 + 2c2 = 6 and 42c3 + 6c2 = 24 how do i get c3 and c2?

OpenStudy (anonymous):

ok first equation change to 2c2=6-12c3 then c2=3-6c3 now substitute into the other equation... 42c3+6(3-6c3)=24... 42c+18-36c=24 6c=6 c3=1

OpenStudy (anonymous):

i changed c3 to c in the middle for my convinience...

OpenStudy (anonymous):

convenience... so you did (didnt matter, i knew ewhat you meant)

OpenStudy (anonymous):

and c2 would be -3?

OpenStudy (anonymous):

haha sorry about my spelling, i know it bugs you. :) and thank you for not making me do the hard parts of the equation!

OpenStudy (anonymous):

sounds right to me!

OpenStudy (anonymous):

how do they get c0 at the end...?

OpenStudy (anonymous):

go back to the t-chart and work backwards. that's what i did.

OpenStudy (anonymous):

hm turns out i didn't need to spend my time on that question after all - it wasn't even taught hduring semester.

OpenStudy (anonymous):

Paul, you haven't been on in a while... Your exam is sometime this week right? Hope it went well :) Or... i hope you do well!